# A solution of the problem of standard compact Clifford-Klein forms

### v1

@@ 2.1 Properties of Lie algebra embeddings yielding compact Clifford-Klein forms
Under the assumptions of Theorem 4, we may assume that
@@
We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the real root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, corresponding to the above decompositions.
@@
$\Delta_{m}=\{\alpha_{c}\in\Delta\,|\,\alpha_{c}|_{\mathfrak{a}}=0\},$
@@
$\Delta_{p}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)>0\},\,\Delta_{n}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)<0\},$
$\Delta_{p}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)>0\},\,\Delta_{n}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)<0\},$
$\Delta_{0}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)=0\},\ \Delta_{0}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)=0\},$
@@
$Z=\sum_{\alpha_{c}\in\Delta_{p}^{+}\cup\Delta_{n}^{+}}\mathfrak{g}_{\alpha_{c}}\subset\mathfrak{n}^{c}.$
Let $\pi:\mathfrak{n}^{c}=\sum_{\alpha_{c}\in\Delta_{m}^{+}}\mathfrak{g}_{\alpha_{c}}\rightarrow Z$ be the natural projection, $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$. Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple and that there exists a non-zero $X\in\mathfrak{t}$ such that $X\not\in(\mathfrak{h}+\mathfrak{l})$. We may assume that there exists a non-zero $X\in\mathfrak{t}$ such that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}.$ Then:
@@
    $Q_{j}\in\sum_{\alpha_{c}\in\Delta_{0},\ \alpha_{c}|_{\mathfrak{a}_{h}}=\gamma}\mathfrak{g}_{\alpha_{c}}.$
@@
We consider triples $(G,H,L)$ which yield Clifford-Klein forms up to an equivalence which is given by the following well-known fact.
Proposition 2.
If a triple $(G,H,L)$ yields a standard compact Clifford-Klein form then so does $(G,g_{1}Hg_{1}^{-1},g_{2}Lg_{2}^{-1})$ for any $g_{1},g_{2}\in G.$
This enables us to choose embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, $\mathfrak{l}\hookrightarrow\mathfrak{g}$ up to conjugacy in $G$, and we will do this throughout without further notice. It is sufficient to consider the condition $G=HL$ on the Lie algebra level, because of the classical result of Onishchik.
Theorem 5 ([29], Theorem 3.1).
Let $\mathfrak{g}$ be a reductive real Lie algebra and let $\mathfrak{g}^{\prime},$ $\mathfrak{g}^{\prime\prime}$ be two reductive subalgebras. Let $G$ be a connected Lie subgroup with Lie algebra $\mathfrak{g}$ and let $G^{\prime}$ and $G^{\prime\prime}$ be its connected Lie subgroups corresponding to Lie algebras $\mathfrak{g}^{\prime}$ and $\mathfrak{g}^{\prime\prime}$. Then $G=G^{\prime}G^{\prime\prime}$ if and only if $\mathfrak{g}=\mathfrak{g}^{\prime}+\mathfrak{g}^{\prime\prime}.$
Theorem 6 ([2], Corollary 3, see also [21]).
If $G/H$ admits a compact Clifford-Klein form then the center of $H$ is compact.
Let
@@
By [2] we may assume that the centers of $H$ and $L$ are compact. Let $\mathfrak{h}^{\prime}\subset\mathfrak{h}$ be a semisimple Lie algebra given by the sum of all simple ideals of $\mathfrak{h}$ which are of non-compact type. Let $H^{\prime}\subset H$ be the closed connected subgroup corresponding to $\mathfrak{h}^{\prime}.$ We see that $L$ acts properly on $G/H^{\prime}$ and that $d(G)=d(H^{\prime})+d(L).$ This way the triple $(G,H^{\prime},L)$ induces a standard compact Clifford-Klein form.
2.2 Chevalley bases compatible with non-compact real forms
Given a root $\alpha_{c}$ and a root space $\mathfrak{g}_{\alpha_{c}}$ we denote by $t_{\alpha}\in\mathfrak{j}^{c}$ the vector determined by the equality $B(t_{\alpha},h)=\alpha_{c}(h)$ for all $h\in\mathfrak{j}^{c}$. Set $h_{\alpha}=2t_{\alpha}/B(t_{\alpha},t_{\alpha}).$
@@ 2.2 Chevalley bases compatible with non-compact real forms
A Chevalley basis of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}$ is a basis of $\mathfrak{g}^{c}$ consisting of $x_{\alpha}\in\mathfrak{g}_{\alpha_{c}}$ and $h_{\alpha}$ with the following properties:
@@
    $[x_{\alpha},x_{-\alpha}]=-h_{\alpha},\,\forall\alpha_{c}\in\Delta$,
@@
    for each pair $\alpha_{c},\beta_{c}\in\Delta$ such that $\alpha_{c}+\beta_{c}\in\Delta$ the constants $c_{\alpha,\beta}\in\mathbb{C}$ determined by $[x_{\alpha},x_{\beta}]=c_{\alpha,\beta}x_{\alpha+\beta}$ satisfy $c_{\alpha,\beta}=c_{-\alpha,-\beta}$.
@@
Lemma 2.
For the chosen $\mathfrak{j}^{c}$ in $\mathfrak{g}^{c}$ there is a base of $\sum_{\alpha_{c}\in\Delta}\mathfrak{g}_{\alpha_{c}}$ of the form
$\{\tilde{x}_{\alpha}\ |\ \tilde{x}_{\alpha}\in\mathfrak{g}_{\alpha_{c}},\ \alpha_{c}\in\Delta\},$
such that for every $\alpha_{c}\in\Delta$
@@
Proof.
Note that $\tau(\mathfrak{j}^{c})=\mathfrak{j}^{c}$ and therefore $\tau$ permutes the root spaces. Since $x_{\alpha}+\tau(x_{\alpha})\in\mathfrak{g}_{u}$ we get $B(x_{\alpha}+\tau(x_{\alpha}),x_{\alpha}+\tau(x_{\alpha}))<0$. Note that $B(x_{\alpha},x_{\beta})\not=0$ if and only if $\alpha_{c}=-\beta_{c}$. Hence $B(x_{\alpha},\tau(x_{\alpha}))<0$ and therefore $\tau(x_{\alpha})=ax_{-\alpha}$ for some $a>0$. Since $\tau$ is an involution, $\tau(x_{-\alpha})=\frac{1}{a}x_{\alpha}$. Therefore, one can take $\tilde{x}_{\alpha}=\frac{1}{\sqrt{ab}}x_{\alpha}$ and $\tilde{x}_{-\alpha}=\frac{\sqrt{a}}{\sqrt{b}}x_{-\alpha}$. ∎
Remark 1.
Note that in this subsection we consider only root systems of semisimple complex Lie algebras, and, therefore, follow the standard notation. However, in the subsequent sections we distinguish between real and complex roots.
@@ 3 Some properties of root systems
In this subsection we establish some properties of abstract root systems. Our aim is Lemma 5 which describes subsets of the root system consisting of roots simultaneously vanishing on a pair of vectors. This may be known, but we did not find any appropriate reference. Let $\Delta\subset\mathbb{R}^{n}$ be an indecomposable root system in the Euclidean space $(\mathbb{R}^{n},(,))$ ([30], Chapter 3). The system $\Delta$ is not assumed to be reduced. Let $\Delta^{+}\subset\Delta$ be a subset of positive roots and $\Pi\subset\Delta^{+}$ the subset of simple roots. There is a unique maximal root $\beta\in\Delta^{+}$ such that for every $\alpha\in\Delta^{+}$ vector $\beta-\alpha$ is a combination of simple roots with non-negative coefficients.
Lemma 3.
Let $A$ be a vector in the interior of the positive Weyl chamber. Then $\beta(A)>\alpha(A)$ for any $\alpha\in\Delta^{+}-\{\beta\}.$
Proof.
Since $A$ is in the interior of the positive Weyl chamber, $\alpha_{i}(A)>0$ for any simple root $\alpha_{i}\in\Pi.$ Thus for $\alpha\in\Delta^{+},$ $\alpha\neq\beta$ we obtain $(\beta-\alpha)(A)>0.$
Proposition 3 ([30], Chapter 3, Proposition 1.2).
Let $\gamma_{1},\gamma_{2}\in\Delta.$ If $(\gamma_{1},\gamma_{2})<0$ then $\gamma_{1}+\gamma_{2}\in\Delta.$
Lemma 4.
If $\beta\in\Delta^{+}$ is such that for any $\alpha\in\Delta^{+}$ the vector $\beta+\alpha$ is not a root then $\beta$ is the highest root.
Proof.
Assume that $\beta,\gamma\in\Delta^{+}$ are distinct and have the property that $\beta+\alpha$ and $\gamma+\alpha$ are not roots for any $\alpha\in\Delta^{+}.$ By Proposition 3 for any simple root $\alpha_{i}$ we get $(\beta,\alpha_{i}),(\gamma,\alpha_{i})\geq 0.$ As $\beta\neq\gamma,$ for some simple root $\alpha_{i}$ we get $(\beta,\alpha_{i})\neq(\gamma,\alpha_{i})$. So $(\beta+\gamma,\alpha_{i})>0.$ By Proposition 3 (setting $\gamma_{1}=\alpha+\beta,$ $\gamma_{2}=-\alpha_{i}$) we see that $\theta:=\beta+\gamma-\alpha_{i}$ is a root. Since $\beta$ is the highest root, $\gamma$ is a positive root and $\alpha_{i}$ is simple (a lowest root) therefore $\gamma-\alpha_{i}=0$ (otherwise $\beta-\theta$ written as a combination of positive roots would have some negative coefficients). But $\gamma$ is a simple root and so does not have the assumed property that $\gamma+\alpha$ is not a root for every positive $\alpha$ (unless $\Delta=\{\alpha,-\alpha\}$ in which case $\gamma=\beta$). ∎
Lemma 5.
Let $X,H\in\mathbb{R}^{n}$ be non-zero. Define
@@
Assume that $\Delta=C_{X}\cup C_{H}.$ Consider sets $C_{H}\setminus C_{X}$ and $C_{X}\setminus C_{H}.$ If one of them is empty then all the roots in $\Delta$ are orthogonal to $X$ or to $H$, but this is impossible. Therefore there exist simple roots $\gamma_{1},\gamma_{2}$ (for any choice of a subset of simple roots) such that
$\gamma_{1}\in C_{H}\setminus C_{X},\ \ \gamma_{2}\in C_{X}\setminus C_{H},$
because any root can be represented as a linear combination of simple roots. Since $C_{H}\cap C_{X}$ is a root system, we can take the set of positive roots $C_{HX}^{+}\subset C_{H}\cap C_{X}.$ Let
$C_{X}^{+}:=\{\alpha\in C_{H}\setminus C_{X}\ |\ \alpha(X)>0\},$
$C_{H}^{+}:=\{\alpha\in C_{X}\setminus C_{H}\ |\ \alpha(H)>0\},$
One can see that for $\alpha\in C_{X}^{+}$ and $\beta\in C_{H}^{+}$, vector $\alpha+\beta$ is not a root, because otherwise one would obtain a root which is non-zero on $H$ and on $X.$ Consider the set
$\tilde{\Delta}^{+}:=C_{HX}^{+}\cup C_{X}^{+}\cup C_{H}^{+}.$
Note that for any $\alpha\in\Delta$ exactly one of the roots $\alpha,$ $-\alpha$ is contained in $\tilde{\Delta}^{+}.$ Also if $\alpha,\gamma\in\tilde{\Delta}^{+}$ and $\alpha+\gamma$ is a root, then $\alpha+\gamma\in\tilde{\Delta}^{+}.$ Therefore $\tilde{\Delta}^{+}$ is the set of positive roots from $\Delta$ with the property
$\forall_{\alpha\in\tilde{\Delta}^{+}}\alpha(X)\geq 0,\ \alpha(H)\geq 0.$
Consider the highest root $\tilde{\beta}\in\tilde{\Delta}^{+}.$ Since $\Delta$ is indecomposable $\tilde{\beta}$ is a linear combination of all simple roots of $\tilde{\Delta}^{+}$ with positive coefficients. But we have seen before that there are simple roots $\tilde{\gamma}_{1},$ $\tilde{\gamma}_{2}\in\tilde{\Delta}^{+}$ such that
$\tilde{\gamma}_{1}(X)>0,\ \ \tilde{\gamma}_{2}(H)>0.$
Thus $\tilde{\beta}(X)>0$ and $\tilde{\beta}(H)>0.$ A contradiction. ∎
Proposition 4 ([6]).
Assume that $X\in\mathfrak{t}$ satisfies the assumptions of Theorem 4 and put $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$. Then
$Z=Z_{h}\oplus[X,Z_{h}]=Z_{l}\oplus[X,Z_{l}]=Z_{h}\oplus Z_{l}=Z_{h}\oplus[X,Z_{l}].$
Lemma 6 ([6]).
Under the assumptions of Theorem 4, The involution $\sigma$ permutes the root spaces given by $\Delta^{+}_{m}$ so that if $\alpha_{c}\in\Delta_{p}^{+}$ then $\sigma(\mathfrak{g}_{\alpha_{c}})\subset\mathfrak{g}_{\alpha_{c}^{{}^{\prime}}}$ for some $\alpha_{c}^{{}^{\prime}}\in\Delta_{n}^{+}$ (by abuse of notation we write $\sigma(\Delta_{p})=\Delta_{n}$). Thus Analogous properties hold for $\Delta_{p}^{-},\Delta_{n}^{-}.$
Lemma 7 ([6]).
None of the non-zero root vectors $x_{\beta}$ for $\beta_{c}\in\Delta_{0}-\Delta_{m}$ centralizes $Z+\tau(Z).$
Proof.
Assume there exists $x_{\beta},\beta_{c}\in\Delta_{0}$ such that $[x_{\beta},Z+\tau(Z)]=0$. Take $H_{\beta}\in i\mathfrak{t}+\mathfrak{a}$, defined, as usual by the condition $B(\tilde{H},H_{\beta})=\beta(\tilde{H})$ for any $\tilde{H}\in i\mathfrak{t}+\mathfrak{a}$. Denote by $\mathfrak{g}^{x_{\beta}}$ the centralizer of $x_{\beta}$ in $\mathfrak{g}^{c}$. By [11], Lemma 3.4.3
@@
$C_{X}=\{\alpha_{c}\in\Delta\,|\alpha_{c}(X)=0\},\,C_{H}=\{\alpha_{c}\in\Delta\,\,|\,\alpha_{c}(H)=0\}.$
@@
Note that for any $\mu_{c}\in\Delta_{m}$, $\mu_{c}(H)=0$. Assume that
@@
but this contradicts Lemma 5. Thus there exists $\alpha_{c},-\alpha_{c}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha_{c}(H)\not=0$. We have one of the following possibilities
@@
    $\alpha_{c}(iT)\neq-\alpha_{c}(H).$ In this case (since one can choose between $\alpha_{c}$ and $-\alpha_{c}$) we may assume that $\alpha_{c}(H_{\beta})<0.$
@@
    $\alpha_{c}(iT)=-\alpha_{c}(H).$ In this case we see that $\sigma(\alpha_{c})(H_{\beta})=\alpha_{c}(-iT+H)=2\alpha_{c}(H)\neq 0$ and by Lemma 6, $\sigma(\alpha_{c})\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ In this case (since one can choose between $\sigma(\alpha_{c})$ and $-\sigma(\alpha_{c})$) we may assume that $\sigma(\alpha_{c})(H_{\beta})<0.$
Therefore in any case there exists $\alpha_{c}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha_{c}(H_{\beta})<0.$ But for any root vector $x_{\alpha}\in Z+\tau(Z)$ we have $\operatorname{ad}_{H_{\beta}}(x_{\alpha})=\alpha_{c}(H_{\beta})x_{\alpha}$ and so for $x_{\alpha}\in\mathfrak{g}^{x_{\beta}},$ $j=\alpha_{c}(H_{\beta})$. This is a contradiction with the inequality $j\geq 0$. The proof of the Lemma is complete. ∎

### v3

@@ 2.2 Properties of Lie algebra embeddings yielding compact Clifford-Klein forms
Let $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ be a standard triple. Then we may assume that
@@
We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the restricted root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, corresponding to the above decompositions.
@@
$\Delta_{m}=\{\alpha\in\Delta\,|\,\alpha|_{\mathfrak{a}}=0\},$
@@
$\Delta_{p}^{+}=\{\alpha\in\Delta_{m}^{+}\,|\,\alpha(iX)>0\},\,\Delta_{n}^{+}=\{\alpha\in\Delta_{m}^{+}\,|\,\alpha(iX)<0\},$
$\Delta_{p}^{-}=\{\alpha\in\Delta_{m}^{-}\,|\,\alpha(iX)>0\},\,\Delta_{n}^{-}=\{\alpha\in\Delta_{m}^{-}\,|\,\alpha(iX)<0\},$
$\Delta_{0}^{+}=\{\alpha\in\Delta_{m}^{+}\,|\,\alpha(iX)=0\},\ \Delta_{0}^{-}=\{\alpha\in\Delta_{m}^{-}\,|\,\alpha(iX)=0\},$
@@
Let us now formulate the basic technical tool used in the proof of the main theorem. This tool is a restriction on the root system of $\mathfrak{g}^{c}$ derived under the assumption that the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ yields a standard compact Clifford-Klein form, but $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$.
@@
Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple of well embedded subalgebras and that $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$. We may assume that there exists a non-zero $X\in\mathfrak{t}$ such that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}$.
@@
$Z=\sum_{\alpha\in\Delta_{p}^{+}\cup\Delta_{n}^{+}}\mathfrak{g}_{\alpha}\subset\mathfrak{n}^{c}.$
Let $\pi:\mathfrak{n}^{c}=\sum_{\alpha\in\Delta_{m}^{+}}\mathfrak{g}_{\alpha}\rightarrow Z$ be the natural projection, $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$. Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple and that there exists a non-zero $X\in\mathfrak{t}$ such that $X\not\in(\mathfrak{h}+\mathfrak{l})$. We may assume that $X\in\mathfrak{t}$ and that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}.$ Then:
@@
    $Q_{j}\in\sum_{\alpha\in\Delta_{0},\ \alpha|_{\mathfrak{a}_{h}}=\gamma}\mathfrak{g}_{\alpha_{c}}.$
@@
Theorem 4 is quite technical, so we think it will be instructive to support it by example which somehow explains the consequences of the relations between complex and real root systems caused by the assumption $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$. These relations together with Lemma 4 yield a contradiction.
Example 1.
We follow notation from Proposition 4, the reader should consult it first. Let $\mathfrak{g}=\mathfrak{so}(n,n+2)$. Looking at Table 4 in [33] one can look at the Satake diagram in this case
[…]
and notice that $\dim\mathfrak{m}_{0}=1$ and that $\Delta_{m}=\emptyset$. It follows easily that necessarily
$X=H_{n+1},\Delta_{p}^{+}=\{\bar{\alpha}_{n}=\alpha_{n}+\alpha_{n+1}\},\Delta_{p}^{-}=\{\bar{\alpha}_{n}=\alpha_{n}-\alpha_{n+1}\}.$
By Lemma 4 every root in $\Delta_{0}-\Delta_{m}$ is a sum of a root from $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and a root from $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}\cup\Delta_{m}$. Since $\Delta_{m}=\emptyset$ we get $\Delta=D_{n+1}$ and therefore it must be
$\Delta=\{\pm\alpha_{n}\pm\alpha_{n+1}\},$
but then the only possibility is $\Delta=D_{2}$, which implies $\mathfrak{g}=\mathfrak{so}(1,3)$ which never yields compact Clifford-Klein forms (with $\mathfrak{h}$ proper and of non-compact type) because the real rank of $\mathfrak{g}$ is 1.
We consider triples $(G,H,L)$ which yield Clifford-Klein forms up to an equivalence which is given by the following well-known fact: if a triple $(G,H,L)$ yields a standard compact Clifford-Klein form then so does $(G,g_{1}Hg_{1}^{-1},g_{2}Lg_{2}^{-1})$ for any $g_{1},g_{2}\in G.$ This enables us to choose embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, $\mathfrak{l}\hookrightarrow\mathfrak{g}$ up to conjugacy in $G$, and we will do this throughout without further notice. By the general theory of compact Clifford-Klein forms and [32], Theorem 3.1 we can work on the Lie algebra level and study Lie algebra decompositions and their relation to the problem of standard compact Clifford-Klein forms.
Also, using the well known result [3] we restrict ourselves to the case when $\mathfrak{h}$ and $\mathfrak{l}$ are semisimple and are direct sums of ideals of non-compact type (one can also consult [8]).
Put
@@
Now we cite a theorem of T. Kobayashi which plays an important role in the proof of the main theorem of this article: it eliminates some (otherwise possible) triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ from the list of candidates.
Theorem 6 ([21], Theorem 1.5).
  • $\mathfrak{a}_{g^{\prime}}\subseteq\mathfrak{a}_{h}$,
  • $d(G^{\prime})>d(H)$.
In this work we consider embeddings of real Lie algebras $\mathfrak{h}\subset\mathfrak{g}$ and their complexifications $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$. Also, if for an embedding of complex Lie algebras $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ there exist real forms $\mathfrak{h}$ of $\mathfrak{h}^{c}$ and $\mathfrak{g}\subset\mathfrak{g}^{c}$ such that $\mathfrak{h}\subset\mathfrak{g}$, we say that the embedding $\mathfrak{h}\subset\mathfrak{g}$ is the real form of the complex embedding $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$.
Given a root $\alpha$ and a root space $\mathfrak{g}_{\alpha}$ we denote by $t_{\alpha}\in\mathfrak{j}^{c}$ the vector determined by the equality $B(t_{\alpha},h)=\alpha(h)$ for all $h\in\mathfrak{j}^{c}$. Set $h_{\alpha}=2t_{\alpha}/B(t_{\alpha},t_{\alpha}).$
@@
A Chevalley basis of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}$ is a basis of $\mathfrak{g}^{c}$ consisting of $x_{\alpha}\in\mathfrak{g}_{\alpha}$ and $h_{\alpha}$ with the following properties:
@@
    $[x_{\alpha},x_{-\alpha}]=-h_{\alpha},\,\forall\alpha\in\Delta$,
@@
    for each pair $\alpha,\beta\in\Delta$ such that $\alpha+\beta\in\Delta$ the constants $c_{\alpha,\beta}\in\mathbb{C}$ determined by $[x_{\alpha},x_{\beta}]=c_{\alpha,\beta}x_{\alpha+\beta}$ satisfy $c_{\alpha,\beta}=c_{-\alpha,-\beta}$.
@@
Lemma 2 ([8]).
For the chosen $\mathfrak{j}^{c}$ in $\mathfrak{g}^{c}$ there is a base of $\sum_{\alpha\in\Delta}\mathfrak{g}_{\alpha}$ of the form
$\{\tilde{x}_{\alpha}\ |\ \tilde{x}_{\alpha}\in\mathfrak{g}_{\alpha},\ \alpha\in\Delta\},$
such that for every $\alpha\in\Delta$
@@ 3 Some properties of root systems
In this subsection we present Lemma 3 (proved in [8]) which describes subsets of the root system consisting of roots simultaneously vanishing on a pair of vectors. Lemma 3 is used in the proof of Lemma 4 which will be essential in the proof of the main result. Let $\Delta\subset\mathbb{R}^{n}$ be an indecomposable root system in the Euclidean space $(\mathbb{R}^{n},(,))$ ([33], Chapter 3). The system $\Delta$ is not assumed to be reduced. Let $\Delta^{+}\subset\Delta$ be a subset of positive roots and $\Pi\subset\Delta^{+}$ the subset of simple roots. There is a unique maximal root $\beta\in\Delta^{+}$ such that for every $\alpha\in\Delta^{+}$ vector $\beta-\alpha$ is a combination of simple roots with non-negative coefficients.
Lemma 3 ([8]).
Let $\Delta\subset\mathbb{R}^{n}$ be an indecomposable root system and $X,H\in\mathbb{R}^{n}-\{0\}$ be non-zero vectors. Define
@@
The following Lemma was proved in [8], but it is so essential, that we repeat the proof here.
Lemma 4.
None of the non-zero root vectors $x_{\beta}$ for $\beta\in\Delta_{0}-\Delta_{m}$ centralizes $Z+\tau(Z).$
@@
Assume there exists $x_{\beta},\beta\in\Delta_{0}$ such that $[x_{\beta},Z+\tau(Z)]=0$. Take $H_{\beta}\in i\mathfrak{t}+\mathfrak{a}$, defined, as usual by the condition $B(\tilde{H},H_{\beta})=\beta(\tilde{H})$ for any $\tilde{H}\in i\mathfrak{t}+\mathfrak{a}$. Denote by $\mathfrak{g}^{x_{\beta}}$ the centralizer of $x_{\beta}$ in $\mathfrak{g}^{c}$. By [12], Lemma 3.4.3
@@
$C_{X}=\{\alpha\in\Delta\,|\alpha(X)=0\},\,C_{H}=\{\alpha\in\Delta\,\,|\,\alpha(H)=0\}.$
@@
Note that for any $\mu\in\Delta_{m}$, $\mu(H)=0$. Assume that
@@
but this contradicts Lemma 3. Thus there exists $\alpha,-\alpha\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha(H)\not=0$. We have one of the following possibilities
@@
    $\alpha(iT)\neq-\alpha(H).$ In this case (since one can choose between $\alpha$ and $-\alpha$) we may assume that $\alpha(H_{\beta})<0.$
@@
    $\alpha(iT)=-\alpha(H).$ In this case we see that $\sigma(\alpha)(H_{\beta})=\alpha(-iT+H)=2\alpha(H)\neq 0$ and by Lemma LABEL:sigma, $\sigma(\alpha)\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ In this case (since one can choose between $\sigma(\alpha)$ and $-\sigma(\alpha)$) we may assume that $\sigma(\alpha)(H_{\beta})<0.$
Therefore in any case there exists $\alpha\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha(H_{\beta})<0.$ But for any root vector $x_{\alpha}\in Z+\tau(Z)$ we have $\operatorname{ad}_{H_{\beta}}(x_{\alpha})=\alpha(H_{\beta})x_{\alpha}$ and so for $x_{\alpha}\in\mathfrak{g}^{x_{\beta}},$ $j=\alpha(H_{\beta})$. This is a contradiction with the inequality $j\geq 0$. The proof of the Lemma is complete. Note that Lemma 2 is tacitly used here. ∎